OpenStudy (anonymous):

find a cubic polynomial equation that has the given roots of -4 and 4i

jimthompson5910 (jim_thompson5910):

Hint: if -4 is a root, then x+4 is a factor if 4i is a root, so is -4i which means x=+-4i x^2 = (+-4i)^2 x^2 = 16i^2 x^2 = -16 x^2+16=0 So if 4i is a root, then x^2 + 16 is a factor